Northern Express paid $85 in dividends and $110 in interest expense during a given year. During that same year, the firm issued $40 in new equity shares, issued new debt of $65, and repaid $23 of old debt. What is the cash flow to creditors for that year?
Answer #1

Cash flow to creditors is calculated as follows:

Cash flow to creditors for a period = Interest expense - Ending long term debt + Beginning long term debt

Thus, according to the information provided, Interest expense = $110

Beginning long term debt = x ( as beginning long term debt is not given )

Ending long term debt = x + 65 - 23

Therefore,

Cash flow to creditors = 110 - (x +65- 23) + x

= 110 - x - 65 + 23 + x

= $ 68